OpenStudy (anonymous):

Which expression will always produce an odd integer if n is any integer? A. N + 1 B.2n C.2(n+3) D.2n - 1

OpenStudy (anonymous):

Please explain how

OpenStudy (kyleighg):

Use the process of elimination and try to replace all of them with several numbers until one fits the requirements

OpenStudy (mudkillz):

It's not a, since if n is 1 you would get 2

OpenStudy (mudkillz):

It's not b, since 2 times 3 is 6

OpenStudy (mudkillz):

It's not c, since 2(1 +3) = 8

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (mudkillz):

So it's d, through process of elimination. 2*1-1=1 2*2-1=3 2*3-1=4 And so on.

OpenStudy (mudkillz):

Can u medal me? Maybe fan 2?